# The cell as a token: high-dimensional geometry in language models and cell embeddings

### v1

@@ Abstract
An emerging area in large-scale pretrained models is inference-time computing, in which large models spontaneously solve new tasks without additional training [51]. For example, large language models can be prompted to produce poetry with meter and scansion that are unseen in their training corpus [89]. The underlying mechanism, in-context learning, exploits the emergent ability of large models not only to retrieve, but also process, information during inference. Achieving such effects within an interpretable embedding space represents a frontier for single-cell foundation models. Several recent models exhibit forms of inference-time reasoning, such as zero-shot embedding of novel cell types, prediction of protein interactions, and anticipation of responses to genetic perturbations [90, 91]. However, it is difficult to clearly assess the difficulty of these tasks, because many informative biological pathways are highly indirect and nonlinear [92, 93]. For example, the immune system implements complex logical operations through elaborate combinatorial receptor-ligand interactions, phosphorylation cascades, and feedback loops, in order to discriminate self from non-self antigens [94, 95]. Parsing these logical circuits is akin to solving a complex mathematical reasoning problem, requiring models that can effectively process symbolic information. In language models, inference-time symbolic reasoning emerges at sufficient scale, with models recently advancing from solving elementary-school word problems to undergraduate mathematics [96, 97, 98]. As single-cell embedding models grow in scale, their ability to identify increasingly intricate regulatory logic will represent a key test of their generalization.
@@ I Acknowledgments
W.G. thanks the members of the Chan-Zuckerberg theory in biology group. This project has been made possible in part by Grant No. DAF2023-329596 from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation. W.G. is supported by NSF DMS 2436233 and NSF CMMI 2440490, as well as Medici Therapeutics.

### v2

@@ Abstract
Large-scale pretrained models exhibit inference-time computation, in which they spontaneously solve new tasks without additional training [66]. For example, large language models can be prompted to produce poetry with meter and scansion that are unseen in their training corpus [117]. The underlying mechanism, in-context learning, exploits the emergent ability of large models not only to retrieve, but also process, information during inference. Inference-time symbolic reasoning appears to improve with model scale, with language models recently advancing from solving elementary-school word problems to standardized mathematics exams for undergraduates [118, 119, 120]. Achieving similar results for biological datasets represents a frontier for single-cell foundation models. Several recent models exhibit forms of inference-time reasoning, such as zero-shot embedding of novel cell types, prediction of protein interactions, and anticipation of responses to genetic perturbations [48, 49]. However, these tasks have unclear difficulty compared to language modeling tasks like standardized exams, leading to conflicting results regarding the efficacy of current single-cell foundation models [121, 122, 123, 124]. A better test may be the ability of large models to decipher the indirect, multiscale, and highly nonlinear logic of many regulatory circuits. For example, the immune system implements elaborate combinatorial receptor-ligand interactions, phosphorylation cascades, and feedback loops, in order to discriminate self from non-self antigens [125, 126]. Parsing these logical circuits is akin to solving a complex mathematical reasoning problem, requiring models that can effectively process symbolic information.
.7 Conclusion & Limitations of the analogy
Drawing parallels between language models and single-cell embeddings reveals shared principles in how high-dimensional spaces encode structured, context-dependent information. However, the analogy between cells and word tokens has natural limits, presaging potential limitations of foundation models for single-cell genomics.
In natural languages, “context” arises from discrete, ordered sequences, where exact position and co-occurrence of tokens convey meaning [2, 3]. In contrast, a cell’s context arises from a web of spatial relationships, signaling interactions, lineage history, and environmental conditions—most of which are not naturally represented as ordered sequences. No two cells are exact replicates, and their surrounding biochemical and environmental context can never be fully reproduced [7, 127]. Furthermore, unlike words in a corpus, cells cannot be resampled from the same underlying distribution without perturbing the system, limiting the robustness and stationarity assumptions of statistical analogies. A key challenge for emerging virtual cell models will thus be their ability to distill informative context in order to resolve polysemy in cell states while still finding concise representations. Other challenges include integrating diverse experimental contexts without loss of biological specificity, and capturing nonlinear regulatory logic within embedding spaces [44, 46, 43, 48].
In contrast to language, where token context is explicit and uniformly structured, biological context is multiscale, incomplete, and often indirect. Moreover, while neighboring words directly inform language token context, a cell’s relevant “neighbors” may be defined in multiple, potentially conflicting ways (physical proximity, developmental stage, functional similarity). Resolving this ambiguity requires contextual cell embeddings, integrating heterogeneous modalities such as spatial transcriptomics, proteomics, chromatin accessibility, or lineage tracing to derive a unified representation. Truly multimodal foundation models offer a potential solution, by treating auxiliary information—like gene ontologies, medical literature, or known regulatory hierarchies—on an even footing with expression data, thus decoupling modality-specific factors from informative biological variation [128, 53, 54, 49]. However, combining modalities at scale raises technical challenges: batch effects, inconsistent coverage among different modalities, and the difficulty of defining context windows across different samples [129, 60]. Even if relevant auxiliary information is available, its incorporation into embeddings can amplify biases in the experimental design, leading to overfitting to specific tissue types, organisms, or experimental protocols [55, 48]. Identifying such effects will be necessary in future virtual cell models, and represents an area where mechanistic interpretability and low-dimensional manifold discovery may prove informative.
The cell token analogy also breaks down when considering the dynamical nature of biological systems. In languages, dynamic embeddings model variability in token meaning without altering the underlying corpus. Yet in biology, a cell’s “meaning” irreversibly changes over time through differentiation, signaling, senescence, and adaptation [61, 86, 130]. Capturing these processes requires embedding models that are temporally aware, capable of representing continuous trajectories, and robust to sparse or noisy longitudinal data. Moreover, truly contextual embeddings for cells must incorporate causal relationships—distinguishing between correlation and regulatory influence—a level of mechanistic grounding without an obvious equivalence in grammatical rules. Improved benchmarks, which test the ability of foundation models to parse complex and indirect regulatory logic, will help help transform cell representations from descriptive maps into predictive, reasoning-ready representations for biology.
@@ I Acknowledgments
W.G. thanks Ski Krieger and the members of the Chan-Zuckerberg theory in biology group. W.G. thanks Medici Therapeutics.This project has been made possible in part by Grant No. DAF2023-329596 from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation. W.G. is supported by NSF DMS 2436233 and NSF CMMI 2440490. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.